Using the provided information below,


x=33.2 , x2=131.67, y=30.78 , y2=116.52image.png, xy=119.8 , and n=10image.png


Find the sample correlation coefficient rimage.png (4mks)

Find the co-efficient of determination


1
Expert's answer
2022-03-21T00:29:20-0400

Correlation coefficient:

"r=\frac{n\Sigma xy-\Sigma x\Sigma y}{\sqrt{(n\Sigma x^2-(\Sigma x)^2)(n\Sigma y^2-(\Sigma y)^2)}}=
\frac{10*119.8-33.2*30.78}{\sqrt{(10*131.67-33.2^2)(10*116.52-30.78^2)}}=0.8148."

Coefficient of determination:

"r^2=0.8148^2=0.6639."